Consider a horizontal 6-mm –thick, 100- mm long straight fin fabricated from steel (k = 57, W/m. K, = 0.5). The base temperature of the fin is 150 OC, while the quiescent air and the surrounding are at 25C. Assume the fin tip is adiabatic and assume the average fin surface temperature is = 125 C :
   a) Draw a clear schematic for the system.            
   b) Estimate the free convection heat transfer coefficient.
c) Check this assumption of and write your comments.     (please check the value)                          
d) Write the equation and its number to be used to calculate the fin heat transfer rate when the fin tip is adiabatic. Assume W = 50 cm, calculate this rate.
Answer #1

-> Ais 7625 I t = 6mm Itin = 125% + - W=0.5m The stoc L=100mm Parumption: a) Air is quiescent medium b) Surface radiation effapper - Nek/he = tf. 77%0.03 & 8.86 w/mk 0.05 for laver surface nu, - tillk = 0.52 +Rai = 0.52 (5.59 718 7/8 = 7.34 hower = N